OpenStudy (anonymous):

Christa will drive sixty-five miles per hour for the next several hours. If she is already eighty-five miles away from her house, how many miles away from her house will she be after three hours? The expression is eighty-five plus sixty-five times h, where h is the number of hours she drives.

OpenStudy (anonymous):

85 + (65*3) = h 85 + 195 = h h = 280

OpenStudy (charlotte123):

@HARSH123 nice!

OpenStudy (anonymous):

@HARSH123: I think I am gonna need an explanation :P

OpenStudy (anonymous):

In one hour she will have gone 65 miles Fifteen more miles to make 85 miles 85 + (65*3) = h 85 + 195 = h h = 280

OpenStudy (anonymous):

i need to go

OpenStudy (charlotte123):

@HARSH123 Yep :D

OpenStudy (charlotte123):

got it cat?

OpenStudy (anonymous):

Yep :D

OpenStudy (charlotte123):

good :)

OpenStudy (anonymous):

Thanks! :3

OpenStudy (charlotte123):

don't tell anyone cat, but I found this: http://answers.yahoo.com/question/index?qid=20120604191529AAWlLlx LOL

OpenStudy (charlotte123):

oh and this too: http://openstudy.com/study#/updates/51fed37ae4b0cc46c14afb8e

OpenStudy (charlotte123):

that can help verify the answer :)

OpenStudy (anonymous):

WHAT?!

OpenStudy (charlotte123):

LAWL

OpenStudy (charlotte123):

enjoy ur moment of realization